17. In a school, students thought of planting trees in and around the school to reduce air pollution. It was decided that the number of trees, that each section of each class will plant, will be the same as the class, in which they are studying, e.g., a section of class I will plant 1 tree, a section of class II will plant 2 trees and so on till class XII. There are three sections of each class. How many trees will be planted by the students?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
It can be observed that the number of trees planted by the students is in an AP. 1, 2, 3, 4, 5………………..12 First term, a = 1 Common difference, d = 2−1 = 1 Sn = n/2 [2a +(n-1)d] S12 = 12/2 [2(1)+(12-1)(1)] = 6(2+11) = 6(13) = 78 Therefore, number of trees planted by 1 section of the classes = 78 Number of trees planted by 3 sections of the classes = 3×78 = 234 Therefore, 234 trees will be planted by the students.